Find the slope and the y-intercept of the equation y − 3(x − 1) = 0.

Answer by fractalier(6550)

From y − 3(x − 1) = 0, we get it into y = mx + b form...
y - 3x + 3 = 0
y = 3x - 3
so the slope is 3 and the y-intercept is -3.